two junglemaen are standing at the two opposite banks of a river facing each other . one of them starts beating a drum and sound reaches to other one after time t1 seconds he starts. then second one starts beating the drum and now first one hear the sound after time t2 seconds. calculate the velocity of sound relative to air and velocity of wind , if it is blowing from first bank to other bank at right angle to the river flow. 

Grade:

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear viswateja ,

let the distance between the banks is d , then t1 =  d /334 + v ;  334 m /s  is the velocity of sound in vaccum , v is the wind speed.   t2 = d /334-v  .divide the two equations ,  t1/t2 = 334-v /334+v   , we get  v = 334(t2 -t1 )/ t2 +t1.

velocity of sound w.r.t wind in first case  =[ 334(t2-t1)/t2+t1 ] + 334 =    668t2/t2+t1 m/s

in second case ,  334 - [ 334(t2-t1)/t2+t1 ] =  668t1/t2+t1 m/s

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.